• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Subedi

    Log entry time 12:14:22 on August21,2009

    Entry number 285130

    keyword=hall work this morning

    I had a plan to work on both left and right spectrometers to see alignment of different happex triggers (left just T1 and T2, right: T1, T2 and T3). I finished aligning left hrs triggers. There is still some issue with L1A and retiming time relationship. Currently the retiming signal is OR of S0 and S2. Bob is suggesting to just drop S2 in the retiming part. This needs more time. Hall is currently under sweeping. I will come after lunch and finish the remaining work taking controlled access. It might take less than an hour probably.